a. Find the pressure on the roof of Grand Hyatt Hotel Manila (one of the tallest buildings in the Philippines if its height is 1,043 feet. b. Find the pressure on the top of Mount Pulag which is 2,922 meters tall.
1
Expert's answer
2021-10-28T17:42:01-0400

Dependency of the barometric pressure on the altitude is given by the formula


"P=P_0\exp(-\dfrac{Mg}{RT}h)"

"P_0=101.325\ kPa, g-=9.807\ m/s^2,"

"M=0.02896\ kg/mol, T=288.15\ K"

"R=8.3143\ (N\cdot m)/(mol\cdot K)"

a. "h=1043\ ft=1043\cdot0.3048\ m=317.9064\ m"



"P=101.325\ kPa\cdot"

"\exp(-\dfrac{0.02896\ kg/mol\cdot9.807\ m/s^2}{8.3143\ (N\cdot m)/(mol\cdot K)\cdot288.15\ K}\cdot317.9064\ m)"

"=97.775\ kPa"

"97.775\ kPa\approx28.8146\ inches\ Hg"


b. "h=2922\ m"



"P=101.325\ kPa\cdot"

"\exp(-\dfrac{0.02896\ kg/mol\cdot9.807\ m/s^2}{8.3143\ (N\cdot m)/(mol\cdot K)\cdot288.15\ K}\cdot2922\ m)"

"=71.660\ kPa"

"71.660\ kPa\approx21.1613\ inches\ Hg"